OpenStudy (anonymous):

How to get the sum of the sequence i(i-1)/2 from i=1 to i=n.

OpenStudy (anonymous):

\[\sum_{i=1}^n\frac{i(i-1)}{2}\]?

OpenStudy (anonymous):

i would start with \[\frac{1}{2}\sum i^2-\sum i\] and use the formula for the sum of squares and the sum of consecutive integers.

OpenStudy (anonymous):

of course i meant one half time the whole thing

OpenStudy (anonymous):

Thank you for your answer. The actual series, sum of (i-1)(i-2)/2 for i=1 to i=n, when generated, is: 0+0+1+3+6+10+15+ . . . This is given as n(n-1)(n-2)/6. How is this result arrived at?